Circular No. 6082

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

o ANDROMEDAE P. Harmanec, Astronomical Institute, Ondřejov, writes: “A series of

reticon observations (619–672 nm) of the B6 IIIe star o And (the primary of a multiple system), secured at the coudé focus of the Ondřejov 2.0-m telescope on Sept. 22/23, shows a dramatic change of the line spectrum in comparison to the previous spectrum of o And obtained with the same instrumentation on Aug. 3/4. New spectra show a double Hα emission with peaks rising some 2–3 percent above the continuum level and a shell absorption with central intensity 0.31, while the August spectrum has Hα in absorption, with a shell core at central intensity 0.37. The total equivalent width of Hα has decreased from 0.628 to 0.416 nm. Other stellar features seen in both August and September spectra are Si II (multiplet 2) lines at 634.7 and 637.1 nm, He I at 667.8 nm, and as-yet-unidentified broad absorption at ∼ 640.06 nm; the Si II and He I lines are stronger in the new spectra. Difference spectra of the present series (created by subtraction of the first spectrum from the remaining ones) reveal the presence of strong rapid variations in all observed stellar lines, with the exception of the unidentified 640-nm feature. The radial velocities of H and He are ∼ −19 km/s on both occasions, while those of the Si II lines are −20 and −10 km/s from the new spectra; all measurements refer to line wings.”

1993 (243) 1 = (243) IDA I (DACTYL) At the recent meetings of the IAU, the presumed satellite of (243) Ida

mentioned on IAUC 5948 was given the permanent designation (243) Ida I and the name Dactyl.

PERIODIC COMET MACHHOLZ 2 (1994o) Several observers have reported that the primary component (A) of this

comet has become diffuse and much fainter in the past week, while component D has brightened and become much more condensed. Total visual magnitude estimates of component A (cf. IAUC 6075): Sept. 17.13 UT, 7.6 (A. Diepvens, Balen, Belgium, 0.15-m refractor); 24.50, 9.0 (C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector); 25.49, 9.0 (Morris; 4′.4 coma); 25.83, 9.4 (J. Kobayashi, Kumamoto, Japan, 0.41-m reflector). Total visual magnitude estimates of component D (cf. IAUC 6081): Sept. 16.09 UT, 9.7 (L. Szentaskó, Veresegyház, Hungary, 0.33-m reflector); 24.50, 8.8 (Morris); 25.50, 8.6 (Morris; 2′.9 coma); 25.83, 9.0 (Kobayashi).

1994 September 26 Daniel W. E. Green